Find the optimum annual equivalent cost for the defender using the following information:

Current trade-in (re-sale) value=$7,000, decreasing by $2000 from the previous year's value

O&M = $2,000 in year 1, increasing by 30% per year

immediate overhaul required (if kept)=$500

Maximum possible lifespan is 5 additional years of service (from now)

Assume MARR = 10%
Answer #1

Using Excel for ESL analysis

Year Discount factor O&M cost PV (O&M) Cumulative (O&M) Cumulative (O&M) + Initial Cost Salvage value PV (Salvage value) NPV (A/P,10%,n) EUAC
A B C D=C*B E F=E+7500 G H=G*B I=F-H J K = I*J
1 0.9091 2000.00 1818.18 1818.18 9318.18 5000.00 4545.45 4772.73 1.1000 5250
2 0.8264 2600.00 2148.76 3966.94 11466.94 3000.00 2479.34 8987.60 0.5762 5179
3 0.7513 3380.00 2539.44 6506.39 14006.39 1000.00 751.31 13255.07 0.4021 5330
4 0.6830 4394.00 3001.16 9507.55 17007.55 -1000.00 -683.01 17690.56 0.3155 5581
5 0.6209 5712.20 3546.83 13054.37 20554.37 -3000.00 -1862.76 22417.14 0.2638 5914
Discount factor 1/(1+0.10)^n
(A/P,i,n) i((1 + i)^n)/((1 + i)^n-1)

Minimum EUAC = 5179 at 2 yr, ESL = 2yrs

AEC at ESL = 5179

Showing formula in excel

Year Discount factor O&M cost PV (O&M) Cumulative (O&M) Cumulative (O&M) + Initial Cost Salvage value PV (Salvage value) NPV (A/P,10%,n) EUAC
A B C D=C*B E F=E+7500 G H=G*B I=F-H J K = I*J
1 =1/(1.1)^A14 =2000 =C14*B14 =D14 =7500+E14 =7000-2000 =G14*B14 =F14-H14 =0.1*((1 + 0.1)^A14)/((1 + 0.1)^A14-1) =I14*J14
2 =1/(1.1)^A15 =C14*1.3 =C15*B15 =E14+D15 =7500+E15 =G14-2000 =G15*B15 =F15-H15 =0.1*((1 + 0.1)^A15)/((1 + 0.1)^A15-1) =I15*J15
3 =1/(1.1)^A16 =C15*1.3 =C16*B16 =E15+D16 =7500+E16 =G15-2000 =G16*B16 =F16-H16 =0.1*((1 + 0.1)^A16)/((1 + 0.1)^A16-1) =I16*J16
4 =1/(1.1)^A17 =C16*1.3 =C17*B17 =E16+D17 =7500+E17 =G16-2000 =G17*B17 =F17-H17 =0.1*((1 + 0.1)^A17)/((1 + 0.1)^A17-1) =I17*J17
5 =1/(1.1)^A18 =C17*1.3 =C18*B18 =E17+D18 =7500+E18 =G17-2000 =G18*B18 =F18-H18 =0.1*((1 + 0.1)^A18)/((1 + 0.1)^A18-1) =I18*J18
Discount factor 1/(1+0.10)^n
(A/P,i,n) i((1 + i)^n)/((1 + i)^n-1)